OpenStudy (amy2012):

danny left college traveling 9 mph. 3 hours later tyler left traveling in the same direction going 18 mph. how long until tyler meets up with danny?

OpenStudy (jsblumen):

Do you know how to set up a problem like this?

OpenStudy (amy2012):

yes...but the answer I put in is not correct...

OpenStudy (jsblumen):

What was the answer you got?

OpenStudy (amy2012):

6

OpenStudy (jsblumen):

That's the answer I got, 6 hours, maybe it is asking for the answer in minutes?

OpenStudy (amy2012):

it is asking for hours....I just got it...it was 3. had to subtract 3 from 6

OpenStudy (jsblumen):

It must be asking how long after Tyler leaves.

OpenStudy (stamp):

danny and tyler will meet once their distances are equal. recall: d = rt danny's rate is 9 mph, and tyler's rate is 18 mph. danny travels for t hours, and tyler travels for 3 less hours since he leaves 3 hours later, so tyler travels t-3 hours.. danny's rt = tyler's rt 9(t) = 19(t -3) 9t = 19t -57 solve for t